# The defect of a cubic threefold

### 08 Dec 2023 v1

@@ 1. Introduction
Inspired by this observation, we study the defect of singular cubic threefolds $X\subset\mathbb{P}^{4}.$ We consider the case when $X$ has isolated (and thus rational) singularities; this is equivalent to $X$ appearing as a hyperplane section of a smooth cubic fourfold [Bro16, Lemma 4.7]. The possible combinations of isolated singularities that can occur were classified in [Vik23] (see also [Yok02], [All03] for a GIT perspective). The defect $\sigma(X)$ is a global invariant of $X$, depending on both the singularities and their position; and can also be defined as
@@
This allows one to compute $\sigma(X)$ and the cohomology of $X$ in terms of the singularity type of $X$, along with the position. As a by-product, we obtain a classification of $\mathbb{Q}$-factorial cubic threefolds.
@@
In [CML09], the authors introduced the notion of a good line in a cubic threefold, in order to study degenerations of the intermediate Jacobian as $X$ becomes mildly singular (see also [CMGHL21]). The main tool is Mumford’s description of the intermediate Jacobian as a Prym variety: for a line on the cubic, projection exhibits the cubic as a conic bundle over $\mathbb{P}^{2}$, whose discriminant curve is a plane quintic with an associated double cover. For the choice of a good line, the singularities of the quintic are in one-to-one correspondence with those of $X$, and one can study the degeneration of the intermediate Jacobian as a Prym variety of the associated double cover, even in the singular situation [Bea77]. This notion was also used in [LSV17], who introduced the stronger notion of a very good line: for such a line, the associated quintic double cover is étale and has both curves irreducible. We show that such a line exists on a cubic threefold $X$, making the Prym construction possible, only when $\sigma(X)=0$.
@@
This sharpens the results of [CML09] and [LSV17]: there are examples of cubic threefolds with worse than allowable (i.e not semi-stable) that none the less admit a very good line.
@@ Acknowledgements
The authors would like to thank Radu Laza for valuable advice. They would like to thank Evgeny Shinder for useful discussions about $\mathbb{Q}$-factoriality and to thank both him and Ivan Cheltsov for helpful comments on a preliminary version. The second author is supported by Methusalem grant METH/21/03 – long term structural funding of the Flemish Government.
@@ 2.1. The projection method
Define $Q_{q}:=V(f_{2}(x_{0},\dots,x_{3}))\subset\mathbb{P}^{3}$, and $S_{q}:=V(g_{3}(x_{0},\dots,x_{3}))$. We let $C_{q}:=Q_{q}\cap S_{q}$ be the $(2,3)$ complete intersection curve, that parametrises lines in $X$ passing through $q$.
@@
This definition agrees with the usual definition of corank of a critical point of a function in this case. Note that if $q\in X$ is a point of corank 3, then $Q_{q}$ is a double plane and the curve $C_{q}$ is non-reduced.
We will use the following well-known lemma:
@@ 2.2. Local Invariants
[Ste97] We define Du Bois invariants for $p\geq 0,q>0$ as
$b^{p,q}(X,x):=\dim H^{q}(Y,\Omega^{p}_{Y}(\log E)(-E)).$
@@
Let $L$ denote the link of the singularity $(X,x)$. We define link invariants:
$l^{p,q}:=\dim H^{q}(E,\Omega^{p}(\log E)\otimes\mathcal{O}_{E})=\dim\mathrm{Gr}^{p}_{F}H^{p+q}(L),$
where $L$ is the link of the singularity.
Using Poincaré duality for $L$ we see that
@@
The Deligne-Hodge numbers of the link $L$ are given by
$h_{i}^{p,q}(L)=\dim\mathrm{Gr}_{F}^{p}\mathrm{Gr}^{W}_{p+q}H^{i}(L).$
We see that $h_{i}^{p,q}=h_{i}^{q,p},$ and $l^{p,i-p}=\sum_{q}h_{i}^{p,q}$. Further, by semi-purity we have $h_{i}^{p,q}=0$ if $i<3$ and $p+q>3$.
@@
Let $(X,x)$ be a rational isolated local complete intersection singularity. Then $l^{p,q}=0$ except possibly for $l^{1,1}$.
@@
The link of a $3$-dimensional isolated complete intersection singularity is simply connected, so $H^{i}(L)\neq 0$ only for $i\in\{0,2,3,5\}$. It follows in this case that $l^{p,q}=0$ unless $p+q\in\{0,2,3,5\}.$ Since $(X,x)$ is a rational singularity, by [Ste97, Lemma 2] $l^{0,i}=l^{i,0}=0$ for all $i$. ∎
@@ 3.2. $\mathbb{Q}$-factoriality
    If $X$ is a cone over a smooth cubic surface, then $\sigma(X)=6$. Further, $Weil(X)/Cart(X)$ is generated by the planes obtained as cones over the lines on the cubic surface.
  2. (2)
@@
For combinations of $ADE$ singularities, $\sigma(X)$ depends on the singularity type and the geometry of $X$, but can still be determined by studying the curve $C_{q}$. We will focus on the $\mathbb{Q}$-factoriality of maximal combinations of singularities - a combination is maximal in a certain class of combinations of singularities if we cannot obtain this combination of singularities by deforming a cubic hypersurface with another combination of singularities from this class. One can easily extract the results for all combinations using [Vik23].
Remark 3.4.
Upper semicontinuity of cohomology implies that if a cubic threefold has defect 0, then all of its deformations have defect 0.
@@
A cubic threefold with $E_{8}+A_{2}$ or $D_{5}+2A_{2}$ singularities is $\mathbb{Q}$-factorial. A cubic threefold with $E_{7}+2A_{1}$ is $\mathbb{Q}$-factorial if and only if it doesn’t contain a plane. Moreover, these are the maximal combinations containing a $D_{n}$ or $E_{n}$ singularity that can be $\mathbb{Q}$-factorial.
@@ 3.3. Mixed Hodge Structure on $H^{3}(X,\mathbb{Z})$
To completely understand the cohomology of the cubic threefold $X$ with isolated singularities, it remains to understand the mixed Hodge structure on the middle cohomology. Let $\{p_{1},\dots p_{r}\}\in X$ be the isolated singular points of $X$. We denote by
$B:=\sum_{i=1}^{r}b^{1,1}(X,p_{i})\text{ and }L:=\sum_{i=1}^{r}l^{1,1}(M_{p_{i}})$
@@
$\underline{h}^{2,1}(X)$ $=5-B;$
$\underline{h}^{1,2}(X)$ $=5-(L-\sigma)-B.$
@@
The vanishing cohomology $H^{3}_{van}$ has a mixed Hodge structure induced through the isomorphism $H^{3}_{van}\cong\bigoplus H^{3}(M_{p_{i}})$, and we can compute the Hodge numbers using Theorem 2.12. In particular, $H^{3}_{\mathrm{van}}$ has no weight 2 cohomology. It follows that $h^{1,1}(H^{3}(X))=\alpha.$
@@ 4. Planes and rational normal cubic scrolls
Recall that a rational normal cubic scroll is a a smooth connected surface $\Sigma\subset\mathbb{P}^{4}$ of minimal degree with $\text{span}(\Sigma)=\mathbb{P}^{4}$. Such a surface is isomorphic to $\mathbb{P}^{2}$ blown up in a point. It is clear that if $X$ contains a plane or a cubic scroll then $\sigma(X)>0$ we include the proof for completeness.
@@
The case of nodes is essentially contained in [FW89]; we provide a proof using the language of this note for completion. Let $q\in X$ be a node, and project to obtain $C_{q}\subset\mathbb{P}^{3}$. The curve $C_{q}$ lies on a smooth quadric, and the singularities of $C_{q}$ are in one-to-one correspondence with the remaining singularities of $X$. Since $\sigma(X)>0$, the curve $C_{q}$ is reducible by Proposition 3.2. It follows that $C_{q}$ has at least $5$ nodes, thus $X$ must have at least $6$ nodes. If the number of nodes is strictly greater than $6$, the curve $C_{q}$ either contains a line, or a plane conic - both imply the existence of a plane in $X$. If $X$ has exactly $6$ nodes in general position, then $X$ contains a cubic scroll by [HT10, Prop. 23]. If the nodes are not in general position, then four nodes must be contained in a plane contained in $X$ by [Avi18, Lemma 2.10]. ∎
@@
Let $q\in X$ be the worst singularity (i.e $A_{n}$ for $n\geq 2)$; the curve $C_{q}\subset\mathbb{P}^{3}$ lies on the quadric cone $Q_{q}$. The blow up $\widetilde{Q}_{q}$ of $Q_{q}$ at the cone point is isomorphic to $\mathbb{F}_{2}$; let $\widetilde{C}\subset\mathbb{F}_{2}$ be the strict transform of $C_{q}\subset Q_{q}$. Applying Proposition 3.2, the curve $\widetilde{C}$ is reducible. Examining the proofs of [Vik23, Propositions 4.6–4.9], we see that whenever $\widetilde{C}$ is reducible, it contains either the premiage of a ruling $\gamma$ of $Q_{q},$ the preimage of a hyperplane section $\Gamma$ of $Q_{q}$ missing the cone point, or the preimage of two twisted cubics $\Gamma_{1},\Gamma_{2}\subset\mathbb{P}^{3}$.
In the first two cases, $X$ clearly contains a plane; it sufices to consider the latter. We claim that $X$ contains a rational normal cubic scroll. Indeed, $\Gamma_{1},\Gamma_{2}$ are two twisted cubics lying on the cubic surface $S_{q}$. Following the argument of [Dol16, Thm 3.5], the linear systems $|\Gamma_{1}|,|\Gamma_{2}|$ define two determinantal representations of the surface $S_{q}$, each obtained from the other by taking the transpose of the matrix. In an identical way to [Dol16, Thm 3.5], the cubic threefold $X$ is isomorphic to a hypersurface $V(\det(A))$, where $A$ is a $3\times 3$ matrix with linear forms in coordinates on $\mathbb{P}^{4}$. By [HT10, Prop 17], it follows that $X$ contains a rational normal cubic scroll (see also [Has00, Lemma 4.1.1]). ∎
@@ 5. Very good lines
This proves Theorem 1.5: the excluded singularities are those of corank $3$ for which $\sigma(X)=0.$
@@ 5.1. Very Good Lines
We first establish the existence of good lines for a cubic threefold with $\sigma(X)=0$. First, we will use the following lemma, which seems to be well known but we could not locate a reference.
@@
Assume first that at least one singular point $q\in X$ is of corank $1$, and let $C_{q}$ be the corresponding curve of lines through $q$. Since $\sigma(X)=0,$ by Proposition 3.2 the curve $C_{q}$ is irreducible. We claim $F(X)$ is birational to $\mathrm{Sym}^{2}(C_{q})$, and is thus irreducible. Let $l\in F(X)\setminus\mathrm{Sing}(F(X))$, and consider the plane $H:=\langle q,l\rangle\subset\mathbb{P}^{4}$. Note that because $\sigma(X)=0,$ this plane is not contained in $X$, and we intersect $H\cap X$. Since $X$ is singular at $q$, we have $H\cap X=l\cup l_{x}\cup l_{y}$, and $l_{x},l_{y}$ both contain $q$. Thus under the projection $\pi:Bl_{q}X\rightarrow\mathbb{P}^{3}$, both $\pi(l_{x})=x$ and $\pi(l_{y})$ are contained in $C_{q}$. Thus $\phi([l])=(x,y)$ defines a rational map $\phi:F(X)\dashrightarrow\mathrm{Sym}^{2}(C_{q}).$ We construct the rational inverse as follows: let $(x,y)\in\mathrm{Sym}^{2}(C_{q})$ where $x\neq y$. Then there exists lines $l_{x},l_{y}$ parametrised by $x,y\in C_{q}$ passing through the singular point $q;$ we take the residual line of the intersection $\langle l_{x},l_{y}\rangle\cap X.$
@@
We claim that $F(X)$ is birational to $C_{1}\times C_{2},$ and is thus irreducible. Let $[l]\in F(X)\setminus\mathrm{Sing}(F(X)),$ and consider the plane $H:=\langle q,l\rangle\subset\mathbb{P}^{4}$. Again, intersecting $H\cap X=l\cup l_{x}\cup l_{y}$ where $l_{x},l_{y}$ are lines in $X$ passing through $q,$, and are thus parametrised by points $x,y\in C_{q}.$ Note that $\pi(l)$ is a line in $\mathbb{P}^{3}$, and thus intersects each $P_{i}$ in a single point ($\pi(l)$ cannot be contained in $P_{i}$ if it was, by analysing the equation of $X$ it must be a component of the $(2,3)$ intersection curve $C_{q}$).
@@
We construct the rational inverse: let $(x,y)\in C_{1}\times C_{2}$, with $x\neq y$. Then there exists lines $l_{x},l_{y}$ parametrised by $x,y\in C_{1}\cup C_{2}$ passing through the singular point $q$; we take the residual line of the intersection $\langle l_{x},l_{y}\rangle\cap X.$
@@
The Fano variety of lines has been well studied in the case of $\sigma(X)>0;$ in particular, [HT10] describes the components of $F(X)$ explicitly for the cubic threefold containing a rational normal cubic scroll (see also [Dol16]); in particular there are 3 components. The birational geometry of a Fano variety for a cubic threefold containing a plane appears in [Had00, §3.3], and in detail in [Bro23] over fields of characteristic not 2. Again, there are three components.

### 08 Jan 2024 v2

@@ 1. Introduction
Inspired by this observation, we study the defect of singular cubic threefolds $X\subset\mathbb{P}^{4}.$ We consider the case when $X$ has isolated (and thus rational) singularities; this is equivalent to $X$ appearing as a hyperplane section of a smooth cubic fourfold [Bro16, Lemma 4.7]. The possible combinations of isolated singularities that can occur were classified in [Vik23] (see also [Yok02], [All03] for a GIT perspective). The defect $\sigma(X)$ is a global invariant of $X$, depending on both the singularities of $X$ and their position; it can also be defined as
@@
This allows one to compute $\sigma(X)$ and the cohomology of $X$ in terms of the singularities of $X$, along with their relative position. As a by-product, we obtain a classification of $\mathbb{Q}$-factorial cubic threefolds.
@@
In [CML09], the authors introduced the notion of a good line in a cubic threefold, in order to study degenerations of the intermediate Jacobian as $X$ becomes mildly singular (see also [CMGHL21]). The main tool is Mumford’s description of the intermediate Jacobian as a Prym variety: for a line on the cubic, projection exhibits the cubic as a conic bundle over $\mathbb{P}^{2}$, whose discriminant curve is a plane quintic with an associated double cover. For the choice of a good line, the singularities of the quintic are in one-to-one correspondence with those of $X$, and one can study the degeneration of the intermediate Jacobian as a Prym variety of the associated double cover, even in the singular situation [Bea77]. In addition, the stronger notion of a very good line was used in [LSV17]: for such a line, the associated quintic double cover is étale and both of the curves are irreducible. We show that there exists a very good line on a cubic threefold $X$, making the Prym construction possible, only when $\sigma(X)=0$.
@@
This sharpens the results of [CML09] and [LSV17]: there are examples of cubic threefolds with worse than allowable singularities (i.e cubics that are not GIT semi-stable) that none the less admit a very good line.
@@ Acknowledgements
The authors would like to thank Radu Laza for valuable advice and illuminating conversations. They would like to thank Evgeny Shinder for useful discussions about $\mathbb{Q}$-factoriality and to thank both him and Ivan Cheltsov for helpful comments on a preliminary version. They are also grateful to Andreas Höring for suggesting Corollary 3.3. The second author is supported by Methusalem grant METH/21/03 – long term structural funding of the Flemish Government.
@@ 2.1. The projection method
Define $Q_{q}:=V(q_{2}(x_{0},\dots,x_{3}))\subset\mathbb{P}^{3}$, and $S_{q}:=V(g_{3}(x_{0},\dots,x_{3}))$. We let $C_{q}:=Q_{q}\cap S_{q}$ be the $(2,3)$ complete intersection curve, that parametrises lines in $X$ passing through $q$. Note that if $Q_{q}$ is a double plane, then the curve $C_{q}$ is non-reduced.
@@
This definition agrees with the usual definition of corank of a critical point of a function in this case. We will use the following well-known lemma:
@@ 2.2. Local invariants
[Ste97] We define the Du Bois invariants of $(X,x)$ as follows: for $p\geq 0,q>0$ we have
$b^{p,q}=b^{p,q}(X,x):=\dim H^{q}(Y,\Omega^{p}_{Y}(\log E)(-E)).$
@@
Let $L_{x}$ denote the link of the singularity $(X,x)$. We define link invariants as
$l^{p,q}=l^{p,q}(X,x):=\dim H^{q}(E,\Omega^{p}(\log E)\otimes\mathcal{O}_{E})=\dim\mathrm{Gr}^{p}_{F}H^{p+q}(L_{x}).$
Using Poincaré duality for $L_{x}$ we see that
@@
The Deligne-Hodge numbers of the link $L_{x}$ are given by
$h_{i}^{p,q}(L_{x})=\dim\mathrm{Gr}_{F}^{p}\mathrm{Gr}^{W}_{p+q}H^{i}(L_{x}).$
We see that $h_{i}^{p,q}(L_{x})=h_{i}^{q,p}(L_{x}),$ and $l^{p,i-p}=\sum_{q}h_{i}^{p,q}(L_{x})$. Further, by semi-purity we have $h_{i}^{p,q}(L_{x})=0$ if $i<3$ and $p+q>3$.
@@
Let $(X,x)$ be a rational isolated local complete intersection singularity. Then $l^{p,q}=0$ except for possibly $l^{1,1}$.
@@
The link of a $3$-dimensional isolated complete intersection singularity is simply connected, so $H^{i}(L_{x})\neq 0$ only for $i\in\{0,2,3,5\}$. It follows in this case that $l^{p,q}=0$ unless $p+q\in\{0,2,3,5\}.$ Since $(X,x)$ is a rational singularity, by [Ste97, Lemma 2] $l^{0,i}=l^{i,0}=0$ for all $i$. ∎
@@ 3.1. The defect of $X$
Corollary 3.3.
Let $X\subset\mathbb{P}^{4}$ be a cubic threefold with only isolated singularities. Then $\sigma(X)\leq 6$. Further, $\sigma(X)=6$ if and only if $X$ is a cone over a cubic surface, and $\sigma(X)=5$ if and only if $X$ is the Segre cubic.
Proof.
It is clear that if $X$ is a cone over a cubic surface, then $\sigma(X)=6;$ indeed, the generators of the group $\mathrm{Weil}(X)/\mathrm{Cart}(X)$ are the planes obtained as cones over the lines on the cubic surface. Suppose then that $X$ is not a cone: if $q\in X$ is a node, then $Q_{q}$ is smooth and $C_{q}$ has at most 6 components. This is only possible if $C_{q}$ has 9 additional nodes: in other words, $X$ is the 10 nodal cubic i.e. the Segre cubic. If $q\in X$ is instead a singularity of type $A_{n}$ with $n>1$, it follows from the results of [Vik23, Section 4] that $C_{q}$ has at most 4 components.
The curve $C_{q}$ corresponding to a corank 2 singularity is a union of two plane cubics, thus it has at most 6 components, and $\sigma(X)\leq 6-2=4$. The curve $C_{q}$ corresponding to a corank 3 singularity has at most 3 components since $Q_{q}$ is a double plane. ∎
@@ 3.2. $\mathbb{Q}$-factoriality
For combinations of $ADE$ singularities, $\sigma(X)$ depends on the singularity type and the geometry of $X$, but can still be determined by studying the curve $C_{q}$. We will focus on the $\mathbb{Q}$-factoriality of maximal combinations of singularities – recall the following definition:
Definition 3.5.
A combination of singularities $T$ on a cubic hypersurface in $\mathbb{P}^{3}$ is maximal in a certain class $\mathcal{C}$ of combinations of singularities if we cannot obtain the combination $T\in\mathcal{C}$ by deforming a cubic hypersurface with another combination of of singularities of class $\mathcal{C}$ in $\mathbb{P}^{3}$.
Additionally, we note that by Proposition 3.2 the defect is upper semi-continuous in families provided that the corank of the worst singularity of each cubic in the family remains constant. This allows us to consider maximal combinations of singularities in the context of propositions below. One can easily extract the $\mathbb{Q}$-factoriality results for all combinations by using [Vik23].
@@
A cubic threefold with $E_{8}+A_{2}$ or $D_{5}+2A_{2}$ singularities is $\mathbb{Q}$-factorial. A cubic threefold with $E_{7}+2A_{1}$ is $\mathbb{Q}$-factorial if and only if it does not contain a plane. Moreover, these are the maximal combinations containing a $D_{n}$ or $E_{n}$ singularity that can be $\mathbb{Q}$-factorial.
@@ 3.3. Mixed Hodge structure on $H^{3}(X,\mathbb{Z})$
To completely understand the cohomology of the cubic threefold $X$ with isolated singularities, it remains to understand the mixed Hodge structure on the middle cohomology. Let $\{q_{1},\dots q_{r}\}\in X$ be the isolated singular points of $X$. We denote by
$B:=\sum_{i=1}^{r}b^{1,1}(X,q_{i})\text{ and }L:=\sum_{i=1}^{r}l^{1,1}(X,q_{i})$
@@
$\underline{h}^{1,2}(X)$ $=5-B;$
$\underline{h}^{2,1}(X)$ $=5-(L-\sigma)-B.$
@@
The vanishing cohomology $H^{3}_{van}$ has a mixed Hodge structure induced through the isomorphism
$H^{3}_{van}\cong\bigoplus H^{3}(M_{q_{i}}),$
where $M_{q_{i}}$ is the Milnor fiber of the singularity $q_{i}$. We can thus compute the Hodge numbers of $H^{3}_{van}$ by using Theorem 2.12. In particular, $H^{3}_{\mathrm{van}}$ has no weight 2 cohomology. It follows that $h^{1,1}(H^{3}(X))=\alpha.$
@@
Corollary 3.11.
If $X$ is a cubic threefold with isolated singularities, then
$h^{3}(X)=10-\mu_{tot}(X)+\sigma(X),$
where $\mu_{tot}(X)$ is the sum of the Milnor numbers of all the singular points of $X$.
Proof.
The statement follows immediately from Lemma 2.10, Theorem 2.12, and Theorem 3.10 above. For the cone case, it is known that $h^{3}(X)=0$ (see e.g. [Dim86, Theorem 2.1]), $\sigma(X)=6$ by Corollary 3.3 and the Milnor number is 16. ∎
@@ 4. Planes and rational normal cubic scrolls
Recall that a rational normal cubic scroll is a a smooth connected surface $\Sigma\subset\mathbb{P}^{4}$ of minimal degree with $\text{span}(\Sigma)=\mathbb{P}^{4}$. Such a surface is isomorphic to $\mathbb{P}^{2}$ blown up in a point. It is clear that if $X$ contains a plane or a cubic scroll then $\sigma(X)>0$ – we include a proof for completeness.
@@
The case of nodes is essentially contained in [FW89]; we provide a proof using the language of this note for completion. Let $q\in X$ be a node, and project to obtain $C_{q}\subset\mathbb{P}^{3}$. The curve $C_{q}$ lies on a smooth quadric, and the singularities of $C_{q}$ are in one-to-one correspondence with the remaining singularities of $X$. Since $\sigma(X)>0$, the curve $C_{q}$ is reducible by Proposition 3.2. It follows that $C_{q}$ has at least $5$ nodes, thus $X$ must have at least $6$ nodes. If the number of nodes is strictly greater than $6$, the curve $C_{q}$ either contains a line, or a plane conic – both imply the existence of a plane in $X$. If $X$ has exactly $6$ nodes in general position, then $X$ contains a cubic scroll by [HT10, Prop. 23]. If the nodes are not in general position, then four nodes must be contained in a plane contained in $X$ by [Avi18, Lemma 2.10]. ∎
@@
Let $q\in X$ be the worst singularity (i.e $A_{n}$ for $n\geq 2)$; the curve $C_{q}\subset\mathbb{P}^{3}$ lies on the quadric cone $Q_{q}$. The blow up $\widetilde{Q}_{q}$ of $Q_{q}$ at the cone point is isomorphic to $\mathbb{F}_{2}$; let $\widetilde{C}\subset\mathbb{F}_{2}$ be the strict transform of $C_{q}\subset Q_{q}$. Applying Proposition 3.2, the curve $\widetilde{C}$ is reducible. Examining the proofs of [Vik23, Propositions 4.6–4.9], we see that whenever $\widetilde{C}$ is reducible, it contains either the preimage of a ruling $\gamma$ of $Q_{q},$ the preimage of a hyperplane section $\Gamma$ of $Q_{q}$ missing the cone point, or the preimage of two twisted cubics $\Gamma_{1},\Gamma_{2}\subset\mathbb{P}^{3}$.
In the first two cases, $X$ clearly contains a plane; it suffices to consider the latter. We claim that $X$ contains a rational normal cubic scroll. Indeed, $\Gamma_{1},\Gamma_{2}$ are two twisted cubics lying on the cubic surface $S_{q}$. Following the argument of [Dol16, Thm 3.5], the linear systems $|\Gamma_{1}|,|\Gamma_{2}|$ define two determinantal representations of the surface $S_{q}$, each obtained from the other by taking the transpose of the matrix. In an identical way to [Dol16, Thm 3.5], the cubic threefold $X$ is isomorphic to a hypersurface $V(\det(A))$, where $A$ is a $3\times 3$ matrix with linear forms in coordinates on $\mathbb{P}^{4}$. By [HT10, Prop 17], it follows that $X$ contains a rational normal cubic scroll (see also [Has00, Lemma 4.1.1]). ∎
@@ 5. Very good lines
This proves Theorem 1.5: the excluded singularities are those of corank $3$ for which $\sigma(X)=0.$ We outline the structure of this section: in §5.1 we will recall the necessary definitions. In particular, we recall the existence of good lines for cubic threefolds with $\sigma(X)=0$. Next in §5.2 we will prove that a cubic containing a plane or a rational normal cubic scroll does not contain a very good line, proving one direction of Theorem 5.1. The remaining direction is proved in §5.3: let us outline the strategy. We prove by contradiction, and suppose that $\sigma(X)=0$ but none of the good lines $l\subset X$ are very good, i.e $\widetilde{D}_{l}$ is reducible. There are two possibilities: either the plane quintic $D_{l}=\widetilde{D}_{l}/\tau$ is reducible or irreducible. We first deal with the case that for all good lines $l\subset X$ the curve $D_{l}$ is reducible, and obtain a contradiction by analysing the singularities of $D_{l}$ (that are in correspondence with that of $X$). We finish the proof by considering the case that for one good line $l$ the curve $D_{l}$ is irreducible: we adapt an argument of [LSV17, §2] to achieve the desired contradiction.
@@ 5.1. Very good lines
In [CML09, §3.2], the authors show that a cubic threefold $X$ has a good line if $\mu(X)\leq 5$, or if the singularities of $X$ are allowable (see also [LSV17, §2]): recall that a cubic threefold $X$ has allowable singularities if it has singularities of type $A_{k}$ with $k\leq 5$ or $D_{4}$ singularities.
We will first extend this result, and prove existence of good lines for a cubic threefold with $\sigma(X)=0$: indeed there exists such cubics with $\sigma(X)=0$ but non-allowable singularities. We will use the following lemma, which seems to be well known but we could not locate a reference.
@@
Assume first that at least one singular point $q\in X$ is of corank $1$, and let $C_{q}$ be the corresponding curve of lines through $q$. Since $\sigma(X)=0,$ by Proposition 3.2 the curve $C_{q}$ is irreducible. We claim $F(X)$ is birational to $\mathrm{Sym}^{2}(C_{q})$, and is thus irreducible. Let $l\in F(X)\setminus\mathrm{Sing}(F(X))$, and consider the plane $H:=\langle q,l\rangle\subset\mathbb{P}^{4}$. Note that because $\sigma(X)=0,$ this plane is not contained in $X$, and we intersect $H\cap X$. Since $X$ is singular at $q$, we have $H\cap X=l\cup l_{x}\cup l_{y}$, and $l_{x},l_{y}$ both contain $q$. Thus under the projection $\pi:Bl_{q}X\rightarrow\mathbb{P}^{3}$, both $\pi(l_{x})=x$ and $\pi(l_{y})$ are contained in $C_{q}$. Thus $\phi([l])=(x,y)$ defines a rational map $\phi:F(X)\dashrightarrow\mathrm{Sym}^{2}(C_{q}).$ We construct the rational inverse as follows: let $(x,y)\in\mathrm{Sym}^{2}(C_{q})$ where $x\neq y$. Then there exists lines $l_{x},l_{y}$ parameterised by $x,y\in C_{q}$ passing through the singular point $q;$ we take the residual line of the intersection $\langle l_{x},l_{y}\rangle\cap X.$
@@
We claim that $F(X)$ is birational to $C_{1}\times C_{2},$ and is thus irreducible. Let $[l]\in F(X)\setminus\mathrm{Sing}(F(X)),$ and consider the plane $H:=\langle q,l\rangle\subset\mathbb{P}^{4}$. Again, intersecting $H\cap X=l\cup l_{x}\cup l_{y}$ where $l_{x},l_{y}$ are lines in $X$ passing through $q,$, and are thus parameterised by points $x,y\in C_{q}.$ Note that $\pi(l)$ is a line in $\mathbb{P}^{3}$, and thus intersects each $P_{i}$ in a single point ($\pi(l)$ cannot be contained in $P_{i}$ if it was, by analysing the equation of $X$ it must be a component of the $(2,3)$ intersection curve $C_{q}$).
@@
We construct the rational inverse: let $(x,y)\in C_{1}\times C_{2}$, with $x\neq y$. Then there exists lines $l_{x},l_{y}$ parameterised by $x,y\in C_{1}\cup C_{2}$ passing through the singular point $q$; we take the residual line of the intersection $\langle l_{x},l_{y}\rangle\cap X.$
@@
The Fano variety of lines has been studied in the case of $\sigma(X)>0.$ For example, in [HT10] the authors describe the components of $F(X)$ explicitly for a general cubic threefold containing a rational normal cubic scroll (see also [Dol16]); in particular there are 3 components. The birational geometry of a Fano variety for a cubic threefold containing a plane appears in [Had00, §3.3], and in detail in [Bro23] over fields of characteristic not 2. Again, there are three components.